Grade 12Physical Chemistry

A balloon having a capacity of 10000 m^3 is filled with helium at 20°c and 1 atm pressure.if the balloons is loaded with 80% of the load that it can lift at ground level at what height will the balloon come to rest ? Assume that the volume of the balloon is constant ,.the atmosphere isothermal,20°c the molecular weight of the air is 28.8 and the ground level pressure is 1 atm the mass of the balloon is 1.3*10^6 g

vaibhav
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
From the Energy Conservation,
Change in potential Energy = Chaane in Kinetic energy.
mgh = P*(V2-V1)
Putting the values we get,
1.3*10^6*10^-3*10*h = 1*(10^7 – 0.8*10^7)
h = (10^7 – 0.8*10^7)/13000 => 2*10^6/13000 => (2/13 )* 10^3 => 0.153*10^3 => 153 meters. approx.